A garment factory produces certain types of output by three machines.The respective daily productions were : Machine A = 4500 units, Machine B =2500 units, Machine C = 3000 units. According to past experience it is found that the fractions of defective output produced by the three machines were 1%,1.2% and 2% RESPECTIVELY.An Item is drawn at random free the day's production run and is found to be defective. What is the probability that it comes from the output of i) Machine A;(ii) Machine B;(iii) Machine C
1
Expert's answer
2019-12-09T11:27:48-0500

Let "D" be the event '' defective item is chosen,'' "A" be the event '' item is produced by machine A'', "B" be the event '' item is produced by machine B'', and "C" be the event '' item is produced by machine C''.


"P(A)={4500 \over 4500+2500+3000}=0.45"

"P(B)={2500 \over 4500+2500+3000}=0.25"

"P(C)={3000 \over 4500+2500+3000}=0.3"

Conditional probabilities


"P(D|A)=0.01, P(D|B)=0.012, P(D|C)=0.02"

i) From Bayes theorem probability of defective item from machine A


"P(A|D)={P(A)P(D|A) \over P(A)P(D|A)+P(B)P(D|AB)+P(C)P(D|C)}="

"={0.45(0.01) \over 0.45(0.01)+0.25(0.012)+0.3(0.02)}\approx0.0296"

ii) Probability of defective item from machine B


"P(B|D)={P(B)P(D|B) \over P(A)P(D|A)+P(B)P(D|AB)+P(C)P(D|C)}="

"={0.25(0.012) \over 0.45(0.01)+0.25(0.012)+0.3(0.02)}\approx0.0222"

iii) Probability of defective item from machine C


"P(C|D)={P(C)P(D|C) \over P(A)P(D|A)+P(B)P(D|AB)+P(C)P(D|C)}="

"={0.3(0.02) \over 0.45(0.01)+0.25(0.012)+0.3(0.02)}\approx0.0444"